The explosion of nitroglycerin, CHS (NO3)3 , is chemically speaking an instantaneous decomposition of a fairly complex compound into several gases, including nitrogen and water vapor. How many grams of nitrogen will result from the explosion of 2900 grams of nitroglycerine ? this equation is already balanced for you! The molar mass of nitro is 227 grams/1 mol

4C3H5(NO3)----12CO2+6N2+O2+10H2O


1
Expert's answer
2021-03-22T06:45:18-0400

Solution:

The balanced chemical equation:

4C3H5(NO3)3 → 12CO2 + 6N2 + O2 + 10H2O

According to the equation above:

Moles of C3H5(NO3)3 / 4 = Moles of N2 / 6

Moles of C3H5(NO3)3 / 2 = Moles of N2 / 3


Moles of C3H5(NO3)3 = Mass of C3H5(NO3)3 / Molar mass of C3H5(NO3)3

The molar mass of C3H5(NO3)3 is 227 g mol-1.

Hence,

Moles of C3H5(NO3)3 = 2900 g / 227 g mol-1 = 12.775 mol


Moles of N2 = 3 × Moles of C3H5(NO3)3 / 2 = (3 × 12.775 mol) / 2 = 19.163 mol


Mass of N2 = Moles of N2 × Molar mass of N2

The molar mass of N2 is 28 g mol-1.

Hence,

Mass of N2 = 19.163 mol × 28 g mol-1 = 536.564 g = 536.56 g

Mass of N2 = 536.56 g


OR (short form solution):

(2900 g C3H5N3O9) × (1 mol C3H5N3O9 / 227 g C3H5N3O9) × (6 mol N2 /4 mol C3H5N3O9) × (28 g N2 / 1 mol N2) = 536.56 g


Answer: 536.56 grams of nitrogen (N2) will result from the explosion of 2900 grams of nitroglycerine.